OpenStudy (anonymous):

Verify the trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. 1. cot x sec^4 x = cot x + 2 tan x + tan^3 x

OpenStudy (irishboy123):

just replace everything by sines and cosines and eventually you will end up with the pythagorean theorem if it is actually an identity.

OpenStudy (anonymous):

i got cot x sec^4 x = cot x sec4x . is that correct?

OpenStudy (danjs):

it just says to match the sides

OpenStudy (anonymous):

oops do i have to include the steps also?

OpenStudy (anonymous):

i know that but i need to know what the steps are

OpenStudy (danjs):

i would just transfer over to only sin and cos, then simplify it, check

OpenStudy (anonymous):

and how do you do that? i have no idea how to transfer over

OpenStudy (danjs):

cot(x) = cos(x) / sin(x) csc = 1/sin .. remember those

OpenStudy (anonymous):

yes

OpenStudy (danjs):

tan = sin/cos

OpenStudy (danjs):

the left side goes to \[\large \frac{ 1 }{ \sin(x)*\cos^3(x) }\]

OpenStudy (danjs):

show the right side does the smae thing

OpenStudy (anonymous):

sec^2(x) = 1 + tan^2(x)?

OpenStudy (danjs):

yeah you can use that on the left side and try to make it look like the right